Please help me solve these two questions.
1. You deposit 3500 in an account that earns 2.5 annual interest. Find the balance after 2 years if the interest is compounded quarterly. Use the formula: A=P (1+r/n)^nt
2. You buy a new car for $35,700. The value of the car decreases by 29% each year. Find the value of the car after 5 years. Use the formula: A=P (1+r/n)^nt


Answer by rothauserc(4718)

Use the given formula A = P(1+r/n)^nt
1) A = 3500(1+.025/4)^4*2 = 3678.88
2) A = 35700(1-.29)^5 = 6441.10